# SMITIN: Self-Monitored Inference-Time INtervention for Generative Music Transformers

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Dynamic control of an audio signal can be considered a fundamental job of most audio professionals. Musicians must precisely control their instruments, and recording studio engineers manipulate various controls to achieve the desired result. Recent advances in the generative modeling of text and images have also been applied to audio signals, leading to an emerging body of literature on how to condition and control audio generative models (Dhariwal et al., 2020; Copet et al., 2023; Liu et al., 2023; Wu et al., 2023a; Garcia et al., 2023; Kreuk et al., 2022; Borsos et al., 2023).
@@
Among possible conditioning inputs, describing the audio a user wishes to generate in terms of natural language arguably provides the most flexibility. Such text-to-music models, however, lack the fine-grained temporal control often desired by audio professionals. This has led to training text-to-music models that allow for additional conditioning inputs, such as melody sequences (Copet et al., 2023), or other types of sequential input (Hawthorne et al., 2022; Donahue et al., 2023), but these models still suffer from the fact that, once trained, the type of conditioning input is fixed. To alleviate this limitation, some works have explored inference-time adaptation approaches that allow the addition of new types of control inputs without requiring retraining, i.e., allowing the users of text-to-music models to build their own controls, or “knobs” (Levy et al., 2023; Wu et al., 2023a; Novack et al., 2024; Lin et al., 2023, 2024). However, much previous work applies mainly to diffusion-based text-to-audio models, or it necessitates training supplementary adapters atop the foundational generative model to introduce additional control. To the best of our knowledge, no one has considered the application of inference-time control techniques to language model-like transformer-based text-to-audio models.
@@
Audio “language models” typically work by first tokenizing chunks of an audio waveform, for example by passing them through the encoder of a pre-trained residual vector quantized autoencoder (Zeghidour et al., 2021; Défossez et al., 2022; Kumar et al., 2023), and then pass these audio tokens through an autoregressive transformer trained to predict the next token. Finally, a decoder converts the generated token into audio. Because this class of text-to-music models shares many architectural characteristics with large language models, it is interesting to explore whether techniques developed to provide inference-time control of LLMs (Subramani et al., 2022; Turner et al., 2023; Hernandez et al., 2023) may also be effective for audio. In particular, we take inspiration from recent work attempting to make text language models more truthful by taking advantage of the fact that the learned internal representations of these models are able to represent the concept of truthfulness (Li et al., 2023). Classifier probes (Alain & Bengio, 2016; Belinkov, 2022; Tenney et al., 2019) applied to each attention head are used to determine whether or not that attention head has learned to represent truthfulness, and inference-time intervention (ITI) is then used to surgically modify the outputs of only the most truth-correlated attention heads, resulting in an LLM that hallucinates less. Previous work (Castellon et al., 2021) has shown that classifier probes trained on the internal representations of generative music transformers can lead to strong performance across a wide variety of music classification tasks. However, these probes were trained using the learned representation from entire layers, not individual attention heads.
@@
We first show using classifier probes that individual self-attention heads in a pre-trained state-of-the-art autoregressive music transformer, MusicGen (Copet et al., 2023), have indeed learned to represent aspects of music we may wish to build control knobs for, making it a priori suitable for ITI approaches such as in (Li et al., 2023). However, we find that direct application of this particular approach to a music transformer is sub-optimal because of the difficulty in generating long temporally coherent music samples, as too much intervention causes the generated music to quickly become incoherent. Furthermore, the method asks for the empirical tuning of the number of heads selected for intervention, impeding the ability to implement multiple intervention kinds at scale. To remedy this, we propose the inclusion of a self-monitoring process into the intervention operation, such that we only apply the intervention when the learned classifier probes tell us it is necessary based on the state of the generation network. Crucially, this self-monitoring technique enables real-time assessment of whether the current generated sample incorporates the target factor, allowing for the generation of musically aligned samples without a costly retraining or fine-tuning process. We also propose to weigh all heads as a function of their learned probe performance, removing the need for empirical head selection tuning.
2 Self-Monitored Inference-Time Intervention
@@ 2.1 Autoregressive Transformer Models
Architectures such as that of MusicGen currently form the foundation of state-of-the-art music generative models. These models are characterized by the autoregressive generation of a sequence of audio frames by transformer models. They include a collection of $L$ multi-head self-attention layers (residual connections, normalization layers, and fully-connected feed-forward layers are employed as usual and not described here). At current time step $t$ in the generated sequence, the $l$-th self-attention layer computes $H$ self-attention heads $z_{l,h}(t)\in\mathbb{R}^{D}$ from an input vector $x_{l}(t)\in\mathbb{R}^{DH}$ as
@@
where $x_{l}(1\!:\!t)=[x_{l}(1),\dots,x_{l}(t)]$, $W^{Q}_{l,h}$, $W^{K}_{l,h}$, and $W^{Q}_{l,h}$ denote the head-specific query, key, and value projection matrices, all in $\mathbb{R}^{D\times DH}$, and Att denotes the attention operator (Vaswani et al., 2017). The output vector $y_{l}(t)\in\mathbb{R}^{DH}$ is obtained by projecting back each head into $DH$-dimensional space and summing:
@@ 2.2 Inference-time Intervention (ITI)
(Li et al., 2023) first recognized that the output of the model could be somewhat controlled at inference by intervening in (i.e., modifying) the computation of the output of the multi-head self-attention layer in Eq. (2). This intervention is done by adding a term to the heads $z_{l,h}(t)\in\mathbb{R}^{D}$ before the projection:
@@
where $\theta_{l,h}\in\mathbb{R}^{D}$ is a vector representing the head-specific ITI “direction” in the $D$-dimensional head-specific space, $\sigma_{l,h}\in\mathbb{R}^{+}$ is a head-specific normalization of $\theta_{l,h}$, $\alpha\in\mathbb{R}^{+}$ is the system-wide ITI strength, and $w_{l,h}\geq 0$ is a head-specific strength weighting. While the formulations are equivalent, we depart somewhat from (Li et al., 2023) by introducing $w_{l,h}$ as a separate term, whereas it was factored into $\theta_{l,h}$ in their notations. This helps to disambiguate the dual function played by $\theta_{l,h}$ in their formulation.
The head-specific ITI directions $\theta_{l,h}$ are obtained through classifier probes. Following (Li et al., 2023), we find a side probing dataset suited to train a classifier mirroring our ITI goal (e.g., drum classifier if the goal is to add drums). We then run the dataset through our (frozen) generation network and collect a set of heads $z_{l,h}(t)$ for all $l,h$. For each $l$ and $h$, we then train a distinct logistic regression classifier probe with parameters $\tilde{\theta}_{l,h}$, whose prediction is obtained as $\text{sigmoid}(\langle\tilde{\theta}_{l,h},z_{l,h}(t)\rangle)$. Once training is complete, we set ITI direction $\theta_{l,h}$ as the final $\tilde{\theta}_{l,h}$, and $\sigma_{l,h}$ as the standard deviation of $\langle\theta_{l,h},z_{l,h}(t)\rangle$ for all $z_{l,h}(t)$ obtained on the combined probing training and testing data. We also take note of the final classifier accuracy $\text{acc}_{l,h}$ on the probing test data. In the case of MusicGen, we apply the generative model in unconditional generation mode for probing, as our probing dataset generally lacks text queries for its audio samples.
In (Li et al., 2023), head-specific weights $w_{l,h}$ are set by finding the top-$K$ heads in terms of classifier probe accuracy $\text{acc}_{l,h}$, setting their $w_{l,h}$ to 1 and the others to 0. The most effective $\alpha$ and $K$ are found by hyperparameter grid search.
@@ 2.4 Soft-weighting
One crucial limitation of (Li et al., 2023) is to leave the number $K$ of heads on which ITI is performed (i.e., the number of pairs $(l,h)$ for which $w_{l,h}(t)\neq 0$) as hyperparameter to tune. We additionally propose a hyperparameter-free soft-weighting approach based on the collected probe accuracies $\text{acc}_{l,h}$, and show that it is sufficient to perform effective ITI (see Tab. 4.1). In practice, we propose setting the weights as
@@ 2.5 Automated Intervention Modulation by Self-monitoring
As our core contribution, we propose to use the classifier probes to drive such a modulation. We first define as $\mathcal{H}_{K}$ the set of top-$K$ heads $(l,h)$ by probe classifier accuracy. For each new generation time step, we collect the set $\mathcal{C}(t)$ of output predictions of the classifier probes for all heads in $\mathcal{H}_{K}$ before intervening on them, i.e.,
@@
      if $\bar{\mathcal{C}}(t)\!<\!\tau$ but $w_{l,h}(t)=0$, we reset to the initial value
@@
3 Experiments
Here, we validate the effectiveness of our proposed contribution by presenting a thorough investigation of ITI applied to the music generative task of instrument addition. This section outlines methods to quantify the effectiveness of ITI techniques objectively, complemented by a subsequent subjective evaluation.
@@ 3.1 Setup
For the evaluation of ITI on music transformer, we utilize the large mono configuration, which is composed of 48 layers with 32 attention heads each, amounting to a total of 1,536 heads. Unless otherwise noted, the self-monitoring configuration of SMITIN is set to an intervention strength of $\alpha=5.0$, a sparse intervention rate of $s=5$, and $K=16$ for the selection of monitoring probes.
To enable the probing mirroring the instrument addition task, we curate probing datasets for drums, bass, guitar, and piano detection from MusDB (Rafii et al., 2019) and MoisesDB (Pereira et al., 2023), both of which provide multi-track recordings with isolated instrument stems. Stems from their training (resp. testing) partitions are used to form the probing training (resp. testing) partitions. To create positive and negative classes, i.e., one including and one excluding the target instrument, we proceed as follows. We collect a first set consisting of the mixed audio from each track recording after removing the target instrument stem, and trim silent sections. We then collect a second set consisting of the target instrument stems, and trim silent sections. We then generate negative samples using audio segments randomly sampled from the first set, and positive samples by randomly mixing a mixture without the target instrument (from the first set) with a stem of the target instrument (from the second set). Following silence trimming, the datasets amount to 8.3, 8.0, 13.1, and 5.7 hours for drums, bass, guitar, and piano, respectively. These durations represent the total amount of paired data available for probing. Since the negative samples are more abundant, we balanced the dataset to ensure an equal number of positive and negative samples, based on the total duration of positive samples.
@@ 3.1.1 Generation Approach
Text-to-Music: This task uses text prompts from the MusicCaps (Agostinelli et al., 2023) dataset to guide the generation process, allowing us to evaluate the effectiveness of ITI in the context of text-conditioned generation. We use the aspect list (e.g., “pop, tinny wide hi hats, mellow piano melody”) over the free text caption to further investigate the impact of adding simple instruction “, add inst.” to the list.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Inferred accuracy of the top-K probes’ monitored decision along the time-axis. The yellow line, green line, and the shaded blue region denote the median, mean, and standard deviation of inferred outputs by the probes, respectively. (Left) Monitored result on a real-world music sample. The high accuracy (close to 1.0) until 3.5 seconds reflects the actual presence of drums, which aligns with the audio sample where drums are present only up to that point. (Right) The four sub-plots display the results of audio continuation on the same input music with varying ITI frequencies ($s=[1,5,10,20]$). These illustrate that more frequent intervention leads to a swifter convergence towards the target musical factor, at the expense of losing musical consistency with the input music.
@@ 3.1.2 Objective Metrics
Fréchet Audio Distance (FAD): FAD (Kilgour et al., 2018) is used to compare generated music with real music datasets, which helps ensure that the intervention does not cause an unrealistic shift in MusicGen’s output distribution. Instead of the conventional approach of using VGGish (Hershey et al., 2017) as the deep encoder, we use L-CLAP mus (Wu et al., 2023b) for dimension reduction, as it has been shown to offer a more accurate representation of music (Gui et al., 2023). As the real music datasets, we adopt MusDB or MoisesDB for audio continuation and MusicCaps for Text-to-Music generation.
@@
Similarity Measurement in Audio Continuation: For the audio continuation task, we measure the musical similarity between input and generated samples. We employ the Music Effects Encoder (Koo et al., 2022), a model trained with self-supervised contrastive learning designed to encode song identity. The encoder is used to embed both input and generated samples, and their cosine similarity is calculated. This metric assesses how well the identity of the input music is preserved in the generated output.
3.2 MusicGen Probing
[…]

### v2

@@ Abstract
{IEEEkeywords}
Classifier probes, music transformer, multi-head attention, music information retrieval.
@@ 1 Introduction
Dynamic control of an audio signal can be considered a fundamental job of most audio professionals. Musicians must precisely control their instruments, and recording studio engineers manipulate various controls to achieve a desired result. Recent advances in the generative modeling of text and images have also been applied to audio signals, leading to an emerging body of literature on how to condition and control audio generative models [1, 2, 3, 4, 5, 6, 7].
@@
Among possible conditioning inputs, describing the audio a user wishes to generate in terms of natural language arguably provides the most flexibility. Such text-to-music models, however, lack the fine-grained temporal control often desired by audio professionals. This has led to training text-to-music models that allow additional conditioning inputs, such as melody sequences [2], or other types of sequential input [8, 9, 10, 11], but these models still suffer from the fact that, once trained, the type of conditioning input is fixed. To alleviate this limitation, some works have explored the addition of new types of control inputs without requiring retraining [12, 4, 13, 14, 15, 16]. The standard approach in most previous work trains supplementary adapters (e.g., LoRA [17]) atop the foundational generative model to introduce additional control, but once the adaptation is complete, it is often not straightforward to vary the strength of the control at inference time. Our goal is to build “knob-like” variable strength controls given only a small number of audio examples both with and without a desired musical trait (i.e., binary labels).
@@
Audio “language models” typically work by first tokenizing chunks of an audio waveform, for example by passing them through the encoder of a pre-trained residual vector quantized autoencoder [18, 19, 20], and then pass these audio tokens through an autoregressive transformer trained to predict the next token. Finally, a decoder converts the generated token back into audio. While not state-of-the-art in terms of generation quality compared to latent diffusion [21, 22] or masked token models [5, 23], autoregressive audio generative models are still useful for many real-time and interactive applications. Furthermore, because this class of text-to-music models shares many architectural characteristics with large language models (LLMs), it is interesting to explore whether techniques developed to provide inference-time control of LLMs [24, 25, 26] may also be effective for audio. In particular, we take inspiration from recent work attempting to make text language models more truthful by taking advantage of the fact that the learned internal representations of these models are able to represent the concept of truthfulness [27]. Classifier probes [28, 29, 30] applied to each attention head are used to determine whether or not that attention head has learned to represent truthfulness, and inference-time intervention (ITI) is then used to surgically modify the outputs of only the most truth-correlated attention heads, resulting in an LLM that hallucinates less. Previous work [31] has shown that classifier probes trained on the internal representations of generative music transformers can lead to strong performance across a wide variety of music classification tasks. However, these probes were trained using the learned representation from entire layers, not individual attention heads.
@@
We first show using classifier probes that individual self-attention heads in a pre-trained autoregressive music transformer, MusicGen [2], have indeed learned to represent aspects of music we may wish to build control knobs for, making it a priori suitable for ITI approaches such as in [27]. However, we find that direct application of this particular approach to a music transformer is sub-optimal because of the difficulty in generating long temporally coherent music samples, as too much intervention causes the generated music to quickly become incoherent. Furthermore, the method asks for the empirical tuning of the number of heads selected for intervention, impeding the ability to implement multiple intervention kinds at scale. To remedy this, we propose the inclusion of a self-monitoring process into the intervention operation, such that we only apply the intervention when the learned classifier probes tell us it is necessary based on the state of the generation network. Crucially, this self-monitoring technique enables real-time assessment of whether the current generated sample incorporates the target factor, allowing for the generation of musically aligned samples without a costly retraining or fine-tuning process. We also propose to weigh all heads as a function of their learned probe performance, removing the need for empirical head selection tuning. The overall idea of the self-monitoring process is outlined in Fig. 1.
The remainder of the paper is organized as follows. In Section 2, we evaluate whether the internal representations from individual attention heads of the pre-trained MusicGen model are suitable for downstream music classification tasks by training simple linear classifier probes. Then, in Section 3, we describe SMITIN, where these classifier probes are used to build custom “knobs” for inference-time control of MusicGen. In Section 4, we assess our proposed approach both objectively and subjectively in terms of intervention success and overall audio quality. Furthermore, we study in Section 5 several important practical aspects of our approach, such as the number of audio examples required to obtain classifier probes sufficient for intervention, whether multiple interventions can be used simultaneously, and the ability of the intervention to prevent the autoregressive generative process from diverging over time into unrealistic and temporally incoherent music. In the Appendix, we additionally provide probing experiments on more downstream tasks and additional details of our subjective and objective results. Our source code will be available online [^1] https://github.com/merlresearch/smitin.
Figure 2: Overview of probing MusicGen. Each audio sample in a labeled dataset is converted to EnCodec tokens and input into MusicGen to predict the next token. The activations for the last time step (orange dots) for each attention head in each layer (blue dots) are used to train a logistic regression classifier (probe).
2 Understanding MusicGen
In this section, we seek to investigate and quantify the comprehension of music by each attention head within MusicGen [2]. MusicGen is a pre-trained and publicly accessible generative music transformer that uses an EnCodec encoder [19] to create discrete audio tokens, an autoregressive transformer to predict the next token, and an EnCodec decoder to output an audio signal. This analysis will provide insights into the model’s potential for fine-grained control via attention-head steering. We use a probing task designed to assess the model’s ability to distinguish musical pieces based on the presence or absence of specific instruments, and explore further downstream tasks in the Appendix.
@@ 2.1 Autoregressive Transformer Models
Architectures such as MusicGen are characterized by the autoregressive generation of a sequence of audio frames by transformer models. They include a collection of $L$ multi-head self-attention layers (residual connections, normalization layers, and fully-connected feed-forward layers are employed as usual and not described here). At current time step $t$ in the generated sequence, the $l$-th self-attention layer computes $H$ self-attention heads $z_{l,h}(t)\in\mathbb{R}^{D}$ from an input vector $x_{l}(t)\in\mathbb{R}^{DH}$ as
@@
where $x_{l}(1\!:\!t)=[x_{l}(1),\dots,x_{l}(t)]$, $W^{Q}_{l,h}$, $W^{K}_{l,h}$, and $W^{Q}_{l,h}$ denote the head-specific query, key, and value projection matrices, all in $\mathbb{R}^{D\times DH}$, and Att denotes the attention operator [32]. The output vector $y_{l}(t)\in\mathbb{R}^{DH}$ is obtained by projecting back each head into $DH$-dimensional space and summing:
@@
2.2 Probing MusicGen
We describe the methodology of probing MusicGen by evaluating the capability of its self-attention heads in recognizing instruments (i.e., determining whether a target instrument is present in the audio stream). The overview of the probing procedure is shown as Fig. 2. We create a dataset by curating data from MUSDB [33] and MoisesDB [34], which offer multi-track recordings with isolated instrument stems. For a given target stem, we form a positive class of mixtures where the target stem is present, and a negative class of corresponding mixtures with the target stem removed as follows. First, we remove from every multi-track recording the time regions where the target stem is silent. Then, out of this pruned recording, the mixture of all of its stems is added to the positive class data, while the mixture of all of its stems except its target stem is added to the negative class data. Subsequently, we process 3-second-long segments of these tracks for training (and testing), passing them through MusicGen to extract the intermediate activation $z_{l,h}(T)$ at the last time step $T$ for every self-attention layer $l$ and head $h$. This forms the basis for the training (and testing) sets of the probe classifier, wherein a simple logistic regression model is employed to distinguish the presence of the instrument.
[…]
3 Self-Monitored Inference-Time Intervention
@@ 3.1 Inference-time Intervention (ITI)
Reference [27] first recognized that the output of the model could be somewhat controlled at inference by intervening in (i.e., modifying) the computation of the output of the multi-head self-attention layer in Eq. (2). This intervention is done by adding a term to the heads $z_{l,h}(t)\in\mathbb{R}^{D}$ before the projection:
@@
where $\theta_{l,h}\in\mathbb{R}^{D}$ is a vector representing the head-specific ITI “direction” in the $D$-dimensional head-specific space, $\sigma_{l,h}\in\mathbb{R}^{+}$ is a head-specific normalization of $\theta_{l,h}$, $\alpha\in\mathbb{R}^{+}$ is the system-wide ITI strength, and $w_{l,h}\geq 0$ is a head-specific strength weighting. While the formulations are equivalent, we depart somewhat from [27] by introducing $w_{l,h}$ as a separate term, whereas it was factored into $\theta_{l,h}$ in their notations. This helps to disambiguate the dual function played by $\theta_{l,h}$ in their formulation.
The head-specific ITI directions $\theta_{l,h}$ are obtained through classifier probes. Following [27], we find an auxiliary probing dataset suited to train a classifier mirroring our ITI goal (e.g., drum classifier if the goal is to add drums). We then run the dataset through our (frozen) generation network and collect a set of heads $z_{l,h}(t)$ for all $l,h$. For each $l$ and $h$, we then train a distinct logistic regression classifier probe with parameters $\tilde{\theta}_{l,h}$, whose prediction is obtained as $\text{sigmoid}(\langle\tilde{\theta}_{l,h},z_{l,h}(t)\rangle)$. Once probe training is complete, we set ITI direction $\theta_{l,h}$ as the final $\tilde{\theta}_{l,h}$, and $\sigma_{l,h}$ as the standard deviation of $\langle\theta_{l,h},z_{l,h}(t)\rangle$ for all $z_{l,h}(t)$ obtained on the combined probing training and testing data. We also take note of the final classifier accuracy $\text{acc}_{l,h}$ on the probing test data. In the case of MusicGen, we apply the generative model in unconditional generation mode for probing, as our probing dataset generally lacks text queries for its audio samples.
In [27], head-specific weights $w_{l,h}$ are set by finding the top-$K$ heads in terms of classifier probe accuracy $\text{acc}_{l,h}$, setting their $w_{l,h}$ to 1 and the others to 0. The most effective $\alpha$ and $K$ are found by hyperparameter grid search.
@@ 3.3 Soft-weighting
One crucial limitation of [27] is that $K$, the number of heads on which ITI is performed (i.e., the number of pairs $(l,h)$ for which $w_{l,h}(t)\neq 0$), is a hyperparameter that must be tuned. We additionally propose a hyperparameter-free soft-weighting approach based on the collected probe accuracies $\text{acc}_{l,h}$, and show that it is sufficient to perform effective ITI (see Appendix Table 6.4). In practice, we propose setting the weights as
@@ 3.4 Automated Intervention Modulation by Self-monitoring
As our core contribution, we propose to use the classifier probes to drive such a modulation. We first define as $\mathcal{H}_{K}$ the set of top-$K$ heads $(l,h)$ by probe classifier accuracy. For each new generation time step $t$, we collect the set $\mathcal{C}(t)$ of output predictions of the classifier probes for all heads in $\mathcal{H}_{K}$ before intervening on them, i.e.,
@@
      if $\bar{\mathcal{C}}(t)\!<\!\tau$ and $w_{l,h}(t)=0$, we reset to the initial value
@@
4 Controlled Experiments on Instrument Addition
Here, we validate the effectiveness of our proposed contribution by presenting a thorough investigation of ITI applied to the music generative task of instrument addition, a task for which evaluation is relatively straightforward. This section outlines methods to quantify the effectiveness of ITI techniques objectively, complemented by a subsequent subjective evaluation.
@@ 4.1 Setup
For the evaluation of ITI on MusicGen, we utilize the large mono configuration, which is composed of 48 layers with 32 attention heads each, amounting to a total of 1,536 heads. Unless otherwise noted, the self-monitoring configuration of SMITIN is set to an intervention strength of $\alpha=5.0$, a sparse intervention rate of $s=5$, and $K=16$ for the selection of monitoring probes.
For the instrument addition task, we curate probing datasets for drums, bass, guitar, and piano detection from MUSDB [33] and MoisesDB [34], both of which provide multi-track recordings with isolated instrument stems. Stems from their training (resp. testing) partitions are used to form the probing training (resp. testing) partition. To create positive and negative classes, i.e., one including and one excluding the target instrument, we proceed as follows. We collect a first set consisting of the mixed audio from each track recording after removing the target instrument stem, and trim silent sections. We then collect a second set consisting of the target instrument stems, and trim silent sections. We then generate negative samples using audio segments randomly sampled from the first set, and positive samples by randomly mixing a mixture without the target instrument (from the first set) with a stem of the target instrument (from the second set). Following silence trimming, the datasets amount to 8.3, 8.0, 13.1, and 5.7 hours for drums, bass, guitar, and piano, respectively. These durations represent the total amount of paired data available for probing. Since the negative samples are more abundant, we balanced the dataset to ensure an equal number of positive and negative samples, based on the total duration of positive samples.
@@ 4.1.1 Generation Approach
Text-to-music: This task uses text prompts from the MusicCaps [35] dataset to guide the generation process, allowing us to evaluate the effectiveness of ITI in the context of text-conditioned generation. We use the aspect list (e.g., “pop, tinny wide hi hats, mellow piano melody”) over the free text caption to further investigate the impact of adding simple instructions, e.g., “add inst.” to the tag list.
@@ 4.1.2 Objective Metrics
Although the probing technique performs well on real-world audio samples, as demonstrated in Section 2, there is some uncertainty regarding its effectiveness on generated samples. Also, because the Success Rate metric uses the same internal monitoring probes of MusicGen as the proposed intervention techniques, there is some concern that it may overestimate success in a way other instrument recognition models would not. Thus, to further validate the Success Rate, we compared its results on all the generated samples used for subjective evaluation in Section 4.4 against decisions from other instrument recognition models. Specifically, we measured the Spearman correlation between the logits produced by our internal probes and those produced by external models such as YAMNet [36], PaSST [37], CLAP [38], and ConvNeXt [39]. YAMNet and PaSST were both trained on AudioSet [40], with PaSST further fine-tuned using the OpenMIC dataset [41]. For CLAP-based recognition, we trained a two-layer MLP to classify the embeddings generated by the frozen CLAP encoder, training on the same dataset as our internal probes. ConvNeXt was trained and tested using the same dataset as well (see Appendix A1 for detailed setup). We used publicly-available weights for YAMNet [^2] https://github.com/tensorflow/models/tree/master/research/audioset/yamnet, PaSST [^3] https://github.com/kkoutini/PaSST, and CLAP [^4] https://github.com/LAION-AI/CLAP.
The results in Table 1 demonstrate a strong positive correlation, confirming that our internal probes align with external models in recognizing the desired musical traits. Furthermore, as we show in the analysis of subjective evaluation in Section 4.4, all the instrument recognition methods also correlate well with human decisions, providing additional validation of the models’ alignment with perceptual recognition.
A key advantage of using the Success Rate metric with MusicGen’s internal probes is that it allows for frame-wise recognition during generation, offering precise detection of musical traits throughout the entire sample. For example, consider a 10-second music sample where drums are only present in the last 3 seconds. External instrument recognition models, which often operate with larger window sizes, may struggle to accurately identify that drums are only present in 30% of the sample. In contrast, our internal probes can precisely distinguish such temporal details, identifying the proportion of the sample that contains the target instrument. This level of granularity is particularly important in controlled generation tasks, where it is critical to accurately capture the presence and distribution of musical traits in the generated audio.
Table 1: Spearman correlation between the Success Rate and other instrument recognition methods. All p-values are below 1e-5.
Correlation Coefficient $\rho$ ($\uparrow$)
Method drums bass guitar piano avg.
YAMNet [36] 0.71 0.50 0.58 0.33 0.53
PaSST [37] 0.69 0.73 0.63 0.40 0.63
CLAP [38] 0.82 0.74 0.68 0.66 0.75
ConvNeXt [39] 0.73 0.84 0.70 0.74 0.75
@@
Fréchet Audio Distance (FAD): FAD [42] is used to compare generated music with real music datasets, which helps ensure that the intervention does not cause an unrealistic shift in MusicGen’s output distribution. Instead of the conventional approach of using VGGish [43] as the deep encoder, we use L-CLAP mus [38] for dimension reduction, as it has been shown to offer a more accurate representation of music [44]. As the real music datasets, we adopt MUSDB (drums/bass) or MoisesDB (guitar/piano) for audio continuation and MusicCaps for text-to-music generation.
@@
Similarity Measurement in audio continuation: For the audio continuation task, we measure the musical similarity between input and generated samples. We employ the Music Effects Encoder [45], a model trained with self-supervised contrastive learning designed to encode song identity. The encoder is used to embed both input and generated samples, and their cosine similarity is calculated. This metric assesses how well the identity of the input music is preserved in the generated output.
4.2 Qualitative Evaluation
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 6: Inferred prediction of the top-K probes’ monitored decision along the time axis. The yellow line, green line, and the shaded blue region denote the median, mean, and standard deviation of inferred outputs by the probes, respectively. The red dashed line indicates threshold value $\tau$ of the current monitoring probes. (Left) Monitored result on a real-world music sample. The high prediction (close to 1.0) until 3.5 seconds reflects the actual presence of drums, which aligns with the audio sample where drums are present only up to that point. (Right) The four sub-plots display the results of audio continuation on the same input music with varying ITI frequencies ($s=[1,5,10,20]$). These illustrate that more frequent intervention leads to a swifter convergence towards the target musical factor, at the expense of losing musical consistency with the input music.
We qualitatively analyze the performance of top-$K$ probes’ monitoring capability through Fig. 6. The red dashed line in each plot indicates the threshold value $\tau$, as mentioned in Section 3.4. From the left side of the figure, it is apparent that the median prediction (yellow line) more closely aligns with the actual presence of the target instrument than the mean prediction (green line). Consequently, we use $\bar{\mathcal{C}}(t)=\text{med}(\mathcal{C}(t))$ to monitor during inference time.
The sub-plots on the right illustrate how the hyper-parameter $s$ influences the outcome of the audio continuation task. We observe a relationship between the intervention frequency $s$ and the rapidity with which the model aligns toward the target factor. This suggests that more frequent interventions cause the transformer to prioritize the target factor more heavily in subsequent token generation, potentially at the expense of referencing the similarity to previous musical context. We discuss this trade-off between intervention quantity and musical coherence throughout this section.
4.3 Objective Evaluation
Utilizing the metrics introduced in Section 4.1.2), we objectively evaluate our methods on audio continuation and text-to-music. We compare with baseline methods that do not incorporate ITI, such as MusicGen’s standard unconditional and text-prompt conditioned generations. Additionally, we benchmark against the original ITI method, which applies a uniform intervention at every time step (i.e., $s=1$) with a configuration of $\alpha=5.0$ and $K=64$ to apply the intervention with the equal amount $\alpha$ to each of the $K$ selected heads. For audio continuation, we examine another ITI approach, weight decay, where $\alpha$ is progressively reduced across time steps during the generation process. For text-to-music, our evaluation includes both our default configuration ($\alpha=5.0$) and a stronger intervention level of $\alpha=10.0$ to observe the trade-off between the intervention strength and the preservation of musical fluency within the generated samples.